Profile Productivity Measurement Objective

In 20x2, Choctaw Company implements a new process affecting labor and materials. The following reported data are provided to evaluate the effect on the company’s productivity:

20x1 20x2
Number of units produced 510,000 360,000
Labor hours used 127,500 72,000
Materials used (lbs.) 1,457,143 900,000

Required:

1. Calculate the productivity profile for 20x1. If required, round your answers to two decimal places.

Partial Operational Productivity Ratios 20x1 Profile
Labor productivity ratio
Material productivity ratio

2. Calculate the productivity profile for 20x2. If required, round your answers to two decimal places.

Partial Operational Productivity Ratios 20x2 Profile
Labor productivity ratio
Material productivity ratio

3. What if the labor hours used in 20x2 were 109,500? What does comparison of the 20x1 and 20x2 profiles now communicate? If required, round your answers to two decimal places.

Partial Operational Productivity Ratios 20x1 Profile 20x2 Profile
Labor productivity ratio
Material productivity ratio
Answer #1

Partial operational productivity ratios

20x1 Profile

1

Labor productivity ratio

Units produced / Labor hours

510,000 / 127,500 = 4

Material productivity ratio

Units produced / Material used

510,000 / 1,457,143 = 0.34

20x2

2

Labor productivity ratio

Units produced / Labor hours

360,000 / 72,000 = 5

Material productivity ratio

Units produced / Material used

360,000 / 900,000 = 0.4

20x1

20x2

3

Labor productivity ratio

Units produced / Labor hours

510,000 / 127,500 = 4

360,000 / 109,500 = 3.28

Material productivity ratio

Units produced / Material used

510,000 / 1,457,143 = 0.34

360,000 / 900,000 = 0.40

Comparison of 20x1 and 20x2:

Labor Productivity ratio:

20x1 = 4 units / hour

20x2 = 3.28 units / hour

Labor Productivity ratio of 20x1 is more than 20x2

Material productivity ratio:

20x1 = 0.34 units / lbs

20x2 = 0.40 units / lbs

Material productivity ratio of 20x2 is more than 20x1.